How many grams of NH


4

Cl

NH4Cl will be formed by this reaction?


1
Expert's answer
2022-04-05T00:53:04-0400

Moles NH3 gas remaining at end of reaction:


Starting moles NH3 = 6.90 g NH3 x 1 mol NH3/17 g = 0.406 moles NH3


6.90 g HCl x 1 mol HCl/36.5 g x 1 mol NH3/mol HCl = 0.189 moles NH3 used up to react with the HCl


moles NH3 remaining = 0.406 - 0.189 = 0.217 moles NH3 remaining


Using the ideal gas law to find the pressure of the gas remaining:


PV = nRT


P = nRT/V = (0.217)(0.0821 Latm/Kmol)(298K)/2.00 L


P = 2.65 atm